# RoboMamba: Multimodal State Space Model for Efficient Robot Reasoning and Manipulation

### v1

@@ Appendix C Additional ablation study
The impact of training strategies on reasoning abilities As shown in Table 4, we explore the impact of different training strategies on common sense reasoning ability. Specifically, we conduct these experiments using input images. In this table, AP refers to alignment pre-training, and IC refers to instruction co-training with different dataset combinations. First, we observe that Ex2 outperforms Ex1 across all three metrics, validating the importance of alignment pre-training, consistent with conclusions drawn by mainstream methods [19, 4, 49]. Next, we find that incorporating the LRV-INSTRUCT 400K dataset (Ex3) indeed improves POPE accuracy, helping our model mitigate the negative effects of hallucinations. Finally, introducing robot-related datasets in co-training not only empowers our model with robot-related reasoning abilities but also enhances common sense reasoning performance, particularly on benchmarks like GQA that are related to spatial reasoning.
| AP | IC(LLaVA-655K) | IC(LRV-400k) | IC(Robo-800k) | OKVQA | GQA | POPE | |
| Ex1 | - | ✓ | - | - | 61.5 | 62.2 | 85.5 |
| Ex2 | ✓ | ✓ | - | - | 62.3 | 62.7 | 85.9 |
| Ex3 | ✓ | ✓ | ✓ | - | 62.0 | 62.6 | 86.6 |
| Ex4 | ✓ | ✓ | ✓ | ✓ | 62.4 | 63.8 | 86.9 |
@@ Appendix D Additional real-world experiments
We conduct real-world experiments involving interactions with various household objects using a Franka Emika robotic arm. We modify the finger gripper by attaching double-sided tape to convert it into a suction gripper, providing the gripper head with adhesive properties. The video demonstrations are included in the supplementary video file. As shown in Figure 5, we visualize our model’s reasoning results on a series of robotic downstream tasks, including long-horizon planning, discriminative affordance, generative affordance, past description, and future prediction. Additionally, we project our RoboMamba’s predicted 3D pose back onto a 2D image based on camera parameter.
@@ Appendix E Reasoning evaluation bencharks
To comprehensively evaluate our model’s capabilities in common sense reasoning, we select several general MLLM evaluation benchmarks, prioritizing those related to robotics. Below is a description of each benchmark
@@
GQA and OCRVQA: These benchmarks are employed to test the model’s ability to identify and comprehend the types and positions of important objects within an image. Such spatial identification skills are vital for tasks related to robotic manipulation and interaction with the environment.
@@
Appendix F Additional related work
Multimodal Large Language Models. Large language models (LLMs) have exhibited remarkable reasoning capabilities across various downstream tasks [19, 90, 2]. When addressing complex multimodal reasoning challenges, multimodal large language models (MLLMs) have shown exceptional visual understanding, i.e., BLIP-2 [49], OpenFlamingo [50], LLaMA-Adapter [19, 51], and LLaVA [52]. Additionally, the introduction of 3D MLLMs [12, 91, 81] seeks to expand the reasoning and conversational capabilities of LLMs to include the 3D modality. However, deploying LMMs is expensive due to their significant computational overhead, primarily caused by their billions of parameters. To mitigate these challenges, recent small-scale models [74, 92] demonstrate impressive performance while maintaining manageable computational costs. LLaVA-Phi [74] empowers the recently developed smaller LLM, Phi-2, for visual instruction tuning. TinyLLaVA [92] and MobileVLM V2 [93] demonstrate that high-quality training data and schemes can effectively compensate for the reasoning abilities of smaller LMMs. Furthermore, Cobra [56] innovatively utilizes an SSM-based language model to reduce complexity. Different from previous works, our goal is to develop an efficient embodied MLLM using the Mamba language model. This model not only possesses common sense understanding but also has the capability to complete manipulation tasks effectively.

### v2

@@ Appendix C Additional ablation study
| Encoder | Image Resolution | OKVQA | GQA | POPE | RoboVQA(BLEU-4) |
| CLIP | 224 x 224 | 46.7 | 50.7 | 79.7 | 35.2 |
| XCiT | 224 x 224 | 63.3 | 64.2 | 86.3 | 42.8 |
| CLIP | 336 x 336 | 62.7 | 63.3 | 87.0 | 41.8 |
| SigLIP | 384 x 384 | 62.4 | 64.4 | 86.0 | 40.6 |
The impact of different image encoders on reasoning abilities In this section, we replace the CLIP encoder used in our initial submission with other linear-complexity encoders, such as XCiT [92]. Additionally, we supplement our experiments by using SigLIP [70] as an image encoder. As shown in Table 4, we analyze the impact of different image encoders and input resolutions on reasoning abilities. The training dataset and strategy remain consistent with those in our main experiment. The results indicate that the choice of image encoder and input resolution does not significantly impact reasoning ability within our RoboMamba VLA framework. However, using an image encoder without cross-modality alignment (i.e., XCiT) presents challenges in converting image tokens to LLM language embeddings. Although our training process includes an alignment pre-training stage, this primarily trains the projection layer. Therefore, in future work, we aim to develop a robotics-specific image encoder capable of projecting image tokens into language embeddings while maintaining linear computational complexity to further improve inference speed.
The impact of training datasets on reasoning abilities As shown in Table 5, we examine the impact of different training datasets on common sense and robotic-specific reasoning abilities. Specifically, we conduct these experiments using input images and the CLIP vision encoder. First, we observe that Ex2 outperforms Ex1 across two benchmarks, confirming that incorporating robotic instruction data can effectively enhance specific reasoning abilities. Similarly, comparing Ex3 and Ex4 shows comparable results, though performance on the GQA benchmark declines. However, in our generated descriptions within robotic scenes, we find that the inclusion of robotic instruction data enhances understanding of geometric relationships. Consequently, we plan to propose a robotic-specific geometric reasoning benchmark to more accurately assess the spatial reasoning capabilities of VLA models. Finally, comparing Ex1, Ex3, and Ex5, we find that using more advanced general instruction datasets does not yield significant performance improvements, which may be due to the model capacity of the 2.7B LLM.
LLaVA 1.5 ShareGPT4V-SFT LLaVA-Next Robo-300k GQA POPE RoboVQA4 Ex1 ✓ - - - 65.3 85.6 26.5 Ex2 ✓ - - ✓ 64.2 86.3 42.8 Ex3 - ✓ - - 64.3 85.2 26.7 Ex4 - ✓ - ✓ 62.1 85.5 42.5 Ex5 - - ✓ - 62.9 86.6 25.1 Ex6 - - ✓ ✓ 60.8 85.4 43.0
@@ Appendix D Additional real-world experiments
We conduct real-world experiments involving interactions with various household objects using a Franka Emika robotic arm. We modify the finger gripper by attaching double-sided tape to convert it into a suction gripper, providing the gripper head with adhesive properties. The video demonstrations are included in the supplementary video file. As shown in Figure 5, we visualize our model’s reasoning results on a series of robotic downstream tasks, including long-horizon planning, discriminative affordance, generative affordance, past description, and future prediction. Additionally, failure cases in reasoning are illustrated in Figure 6. Compared to the ground truth, RoboMamba demonstrates limitations in reasoning ability on some complex tasks, occasionally misinterpreting the current task objective or the target manipulated object.
@@ Appendix E Reasoning evaluation bencharks
GQA: These benchmarks are employed to test the model’s ability to identify and comprehend the types and positions of important objects within an image. Such spatial identification skills are vital for tasks related to robotic manipulation and interaction with the environment.